### v3

# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

Abstract.
[…]
2020 Mathematics Subject Classification. Primary 37B05; Secondary 22E15, 22D45
Keywords. Distal actions of automorphisms, Lie groups, one-parameter subgroups, Chabauty topology.
1. Introduction
For a Hausdorff topological space $X$, a homeomorphism $T$ of $X$ is said to be distal (equivalently, $T$ acts distally on $X$) if for any pair of distinct elements $x,y\in X$, the closure of the double orbit $\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}$ in $X\times X$ does not intersect the diagonal, i.e. for $x,y\in X$ with $x\neq y$, $\overline{\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}}\cap\{(a,a)\mid a\in X\}=\emptyset$. David Hilbert introduced the notion of distality to study non-ergodic actions on compact spaces (see Ellis [16] and Moore [30]). Distal actions on compact spaces, Lie groups as well as on locally compact groups have been studied by many mathematicians in different contexts (see Abels [1, 2], Choudhuri et al [10], Ellis [16], Furstenberg [18], Palit-Shah [31], Palit-Prajapati-Shah [32], Raja-Shah [34, 35], Shah [37], Shah-Yadav [39, 41, 42], and references cited therein).
For a locally compact (Hausdorff) group $G$ with the identity $e$, let ${\rm Sub}_{G}$ denote the set of all closed subgroups of $G$ endowed with the Chabauty topology. Recall that ${\rm Sub}_{G}$ is compact and Hausdorff, and it is metrisable if $G$ is second countable [6]. Let ${\rm Aut}(G)$ denote the group of all automorphisms of $G$; an automorphism is both a homomorphism and a homeomorphism. There is a canonical group action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$; namely, $(T,H)\mapsto T(H)$, $T\in{\rm Aut}(G)$, $H\in{\rm Sub}_{G}$; this action gives a homomorphism from ${\rm Aut}(G)$ to ${\rm Homeo}({\rm Sub}_{G})$. As the image of ${\rm Aut}(G)$ under this action is a large subclass of homeomorphisms of ${\rm Sub}_{G}$, it is important to study the dynamics of the action of this special subclass in terms of distality.
In [42], Shah and Yadav have studied the action of automorphisms of a connected Lie group $G$ on ${\rm Sub}_{G}$ extensively and have shown in particular that for a large class of $G$, which does not have any compact central subgroup of positive dimension, an automorphism $T$ of $G$ acts distally on ${\rm Sub}^{a}_{G}$, the space of closed abelian subgroups of $G$, if and only if $T$ generates a relatively compact subgroup in ${\rm Aut}(G)$. A similar characterisation also holds for $T$ belonging to the class (NC), i.e. those $T$ for which the closure of the $T$-orbit of any discrete cyclic group does not contain the trivial group $\{e\}$. Note that ${\rm Sub}^{a}_{G}$ is very large for many groups $G$; e.g. it is the same as ${\rm Sub}_{G}$ if $G$ is abelian. It also contains all discrete cyclic subgroups. Shah, with Palit in [31], and with Palit and Prajapati in [32], has studied distal actions of automorphisms of a discrete group $G$ on ${\rm Sub}^{a}_{G}$, where $G$ is either polycyclic or a lattice in a connected Lie group.
The question arises whether one can characterise distal actions of automorphisms on other smaller invariant subspaces of ${\rm Sub}^{a}_{G}$ for a connected Lie group $G$. Most of the techniques in earlier result mentioned above involve study of the actions on a class of discrete abelian subgroups. Here, we consider the class of smallest closed connected abelian subgroups; namely, closed one-parameter subgroups of $G$ and study the dynamics of distal actions of automorphisms on it. Let ${\rm Sub}^{p}_{G}$ denote the smallest closed subset of ${\rm Sub}_{G}$ containing all closed one-parameter subgroups of $G$. Note that ${\rm Sub}^{p}_{G}$ is compact and it is invariant under the action of ${\rm Aut}(G)$. We get a characterisation for a large class of $G$ as follows.
Theorem 1.1.
Let $G$ be a connected Lie group without any non-trivial compact connected central subgroup and let $T$ be an automorphism of $G$. Then the following hold:
  1. (1)
    If $G$ is abelian, then $G$ is isomorphic to ${\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$, and $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if $T\in{\mathcal{K}}{\mathcal{D}}$, where ${\mathcal{K}}$ is a compact subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{D}}$ is the center of ${\rm GL}(d,{\mathbb{R}})$.
  2. (2)
    If $G$ is not abelian, then $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if $T\in{\mathcal{K}}$, a compact subgroup of ${\rm Aut}(G)$.
Note that the condition on the center in Theorem 1.1 holds if $G$ is simply connected, or more generally its nilradical is simply connected. It also holds for $G/M$ for any connected Lie group $G$, where $M$ is the maximal central torus in $G$ (the maximal compact connected central subgroup in $G$). The result in the case of non-abelian groups $G$ in (2) as above gives a similar characterisation as in Theorem 4.1 of [42], even though the space ${\rm Sub}^{p}_{G}$ is much smaller than ${\rm Sub}^{a}_{G}$. In case $G$ as above is abelian, then $G={\mathbb{R}}^{d}$, a vector group, and ${\rm Sub}^{p}_{G}$ is homeomorphic to ${\mathbb{RP}}^{d-1}\sqcup\{\{0\}\}$, and the Theorem 1.1 (1) shows that a larger class of $T$ acts distally on ${\rm Sub}^{p}_{G}$.
In [42], it is shown for a connected Lie group $G$ without any compact central subgroup of positive dimension that if $T$ acts distally on ${\rm Sub}^{a}_{G}$, then it acts distally on $G$ (more generally, see Theorem 3.6 and Corollary 3.7 of [42]). We generalise this for Lie groups which are not vector groups; as in the later case, Theorem 1.1 (1) shows that it is possible to have automorphisms which act distally on ${\rm Sub}^{p}_{G}$ but not on $G$.
Theorem 1.2.
Let $G$ be a connected Lie group which is not a vector group and let $T\in{\rm Aut}(G)$ be such that it acts distally on the largest compact connected central subgroup of $G$. If $T$ acts distally on ${\rm Sub}^{p}_{G}$, then $T$ acts distally on $G$.
An automorphism $T$ of a connected Lie group $G$ is said to be unipotent if all the eigenvalues of ${\rm d\,}T$ are equal to 1, where ${\rm d\,}T$ is the corresponding Lie algebra automorphism of the Lie algebra of $G$. It is known that $T$ acts distally on $G$ if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf. [1]). Any unipotent automorphism acts distally on $G$ and if it belongs to a compact subgroup of ${\rm Aut}(G)$, then it is trivial. Theorem 4.3 of [42] shows that no non-trivial unipotent automorphism acts distally on ${\rm Sub}^{a}_{G}$; generalising this statement we prove the following.
Theorem 1.3.
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$ be unipotent. Then $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if $T={\rm Id}$, the identity map.
In particular, Theorem 1.3 illustrates that the converse of Theorem 1.2 does not hold as the non-trivial unipotent automorphisms do not act distally on ${\rm Sub}^{p}_{G}$.
Let ${\rm Sub}^{c}_{G}$ denote the set of all discrete cyclic subgroups of $G$. It is invariant under the action of ${\rm Aut}(G)$ and so is its closure in ${\rm Sub}_{G}$. In [31] and [32], distal actions of automorphisms of ${\rm Sub}^{c}_{\Gamma}$ have been characterised, where $\Gamma$ is a certain type of discrete group, e.g. a discrete polycyclic group or a lattice in a connected Lie group. In a connected Lie group $G$, any closed one-parameter subgroup $H=\{x_{t}\}_{t\in{\mathbb{R}}}$ is the limit of $\{H_{n}\}_{n\in{\mathbb{N}}}$ as $n\to\infty$ in ${\rm Sub}_{G}$, where $H_{n}=\{x_{m/n}\mid m\in{\mathbb{Z}}\}$, $n\in{\mathbb{N}}$, are discrete cyclic groups. Thus ${\rm Sub}^{p}_{G}\subset\overline{{\rm Sub}^{c}_{G}}$ (the closure of ${\rm Sub}^{c}_{G}$ in ${\rm Sub}_{G})$ and the latter is contained in ${\rm Sub}^{a}_{G}$. We should note that $\overline{{\rm Sub}^{c}_{G}}$ is much bigger than ${\rm Sub}^{p}_{G}$ as the former contains all discrete cyclic subgroups (see Lemma 2.2). If $T$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$, then $T\in$ (NC). It is easy to see that Theorems 1.1 (2), 1.2 and 1.3 remain valid if ${\rm Sub}^{p}_{G}$ is replaced by $\overline{{\rm Sub}^{c}_{G}}$ in the respective statements.
The following generalises a part of Theorem 4.1 of [42] and also extends Theorem 1.1 (2).
Theorem 1.4.
Let $G$ be a connected Lie group such that it is not a vector group. Let $T\in{\rm Aut}(G)$ be such that it acts distally on the largest compact connected central subgroup of $G$. Then the following are equivalent:
  1. (1)
    $T$ acts distally on ${\rm Sub}^{p}_{G}$.
  2. (2)
    $T$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$.
  3. (3)
    $T$ acts distally on ${\rm Sub}^{a}_{G}$.
  4. (4)
    $T$ acts distally on ${\rm Sub}_{G}$.
  5. (5)
    $T$ is contained in a compact subgroup of ${\rm Aut}(G)$.
We may note that in case $G$ is a vector group, Theorem 1.1 (1) shows that $(1)$ and $(5)$ above are not equivalent; however, in this case, statements $(2-5)$ above are equivalent by Theorem 4.1 of [42] as (2) above implies that $T\in({\rm NC})$.
Recall that a group $\Gamma$ of homeomorphisms of a (Hausdorff) topological space $X$ acts distally on $X$ if $\overline{\{(\gamma(x),\gamma(y))\mid\gamma\in\Gamma\}}\cap\{(d,d)\mid\ d\in X\}=\emptyset$ for any $x,y\in X$ with $x\neq y$. In particular, a homeomorphism $T$ of $X$ acts distally on $X$ if and only if the cyclic group $\{T^{n}\mid n\in{\mathbb{Z}}\}$ acts distally on $X$. For a Lie group $G$, any compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. Lemma 2.2 in [41]). It follows from Theorem 1.1 of [2] that for a subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$, ${\mathcal{H}}$ acts distally on $G$ if and only if ${\mathcal{H}}$ acts distally on the Lie algebra of $G$, and these statements are equivalent to the statement that every element of ${\mathcal{H}}$ acts distally on $G$ (see also [11]). Let $\overline{{\mathcal{H}}}$ denote the closure of ${\mathcal{H}}$ in ${\rm Aut}(G)$. It is easy to see using Abel’s results in [1, 2] that ${\mathcal{H}}$ acts distally on $G$ if and only if so does $\overline{{\mathcal{H}}}$. It follows from Ellis’ result in [16] that ${\mathcal{H}}$ acts distally on any compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$.
Using Theorems 1.1 and 1.4, we get the following theorem for the action of any subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$. If ${\mathcal{H}}$ is closed, the theorem shows that ${\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if so does every element of ${\mathcal{H}}$.
Theorem 1.5.
Let $G$ be a connected Lie group. Let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$ such that it acts distally on the largest compact connected central subgroup of $G$. Consider the following statements:
  1. (1)
    Every element of $\overline{{\mathcal{H}}}$ acts distally on ${\rm Sub}^{p}_{G}$.
  2. (2)
    ${\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$.
  3. (3)
    ${\mathcal{H}}$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$.
  4. (4)
    ${\mathcal{H}}$ acts distally on ${\rm Sub}^{a}_{G}$.
  5. (5)
    ${\mathcal{H}}$ acts distally on ${\rm Sub}_{G}$.
  6. (6)
    $\overline{{\mathcal{H}}}$ is a compact group.
Then, $(1)$ and $(2)$ are equivalent and $(3-6)$ are equivalent.
If $G$ is not a vector group, then $(1-6)$ are equivalent.
If $G$ is a vector group, i.e. if $G={\mathbb{R}}^{d}$, for some $d\in{\mathbb{N}}$, then $(1)$ and $(2)$ are equivalent to the following:
  1. (7)
    ${\mathcal{H}}\subset{\mathcal{K}}{\mathcal{D}}$, where ${\mathcal{K}}$ is a compact subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{D}}$ is the center of ${\rm GL}(d,{\mathbb{R}})$.
Note that in Theorem 1.5, one can also have equivalent statements similar to $(1)$ and $(2)$, where ${\rm Sub}^{p}_{G}$ is replaced by any of $\overline{{\rm Sub}^{c}_{G}}$, ${\rm Sub}^{a}_{G}$ or ${\rm Sub}_{G}$. We may note here that one can not replace $\overline{{\mathcal{H}}}$ by ${\mathcal{H}}$ in (1) of Theorem 1.5, as there is a subgroup ${\mathcal{H}}$ of ${\rm SL}(4,{\mathbb{C}})$ in which every element of ${\mathcal{H}}$ is contained in a compact subgroup of ${\rm SL}(4,{\mathbb{C}})$ but the closure of ${\mathcal{H}}$ in ${\rm SL}(4,{\mathbb{C}})$ is connected and contains unipotent elements (cf. [5]). The condition that $T$ (resp. a subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$) acts distally on the maximal compact connected central subgroup (maximal central torus) of $G$ is satisfied easily if $T\in{\rm Aut}(G)^{0}$, (resp. ${\mathcal{H}}\subset{\rm Aut}(G)^{0}$), as ${\rm Aut}(G)^{0}$ acts trivially on any compact central subgroup. It is also satisfied if $T^{n}\in{\rm Aut}(G)^{0}$ for some $n\in{\mathbb{N}}$ (resp. ${\mathcal{H}}$ or ${\mathcal{H}}{\rm Aut}(G)^{0}$ has finitely many connected components). Here, ${\rm Aut}(G)^{0}$ denotes the connected component of the identity in ${\rm Aut}(G)$.
Let ${\rm Inn}(G)$ denote the group of inner automorphisms of $G$. It is a (not necessarily closed) normal subgroup of ${\rm Aut}(G)$. The following corollary generalises the first part of Corollary 4.5 in [42].
Corollary 1.6.
Let $G$ be a connected Lie group. Then the following are equivalent:
  1. (1)
    Every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$.
  2. (2)
    Every inner automorphism of $G$ acts distally on ${\rm Sub}_{G}$.
  3. (3)
    ${\rm Inn}(G)$ acts distally on ${\rm Sub}^{p}_{G}$.
  4. (4)
    ${\rm Inn}(G)$ acts distally on ${\rm Sub}_{G}$.
  5. (5)
    $G$ is compact or $G={\mathbb{R}}^{n}\times K$, for some $n\in{\mathbb{N}}$ and the maximal compact normal subgroup $K$ of $G$.
In Corollary 1.6, one can also write equivalent statements involving the action on $\overline{{\rm Sub}^{c}_{G}}$.
Note that ${\rm Sub}^{p}_{G}$ is contained in a larger compact set ${\rm Sub}^{co}_{G}$, which is the closure of the set of all closed connected subgroups of $G$. Hence it is easy to show that Theorems 1.11.3 are all valid if ${\rm Sub}^{p}_{G}$ is replaced by ${\rm Sub}^{co}_{G}$. In Theorems 1.4 and 1.5 and Corollary 1.6, the action of a given automorphism or a subgroup of ${\rm Aut}(G)$ is distal on ${\rm Sub}^{co}_{G}$ if and only if it is so on ${\rm Sub}^{p}_{G}$, under the condition on the group and the action on the central torus.
[…]
We now fix some notations. Let $G$ be a connected Lie group. Let $R$ (resp. $N$) denote the radical (resp. nilradical) of $G$, i.e. the largest connected solvable (resp. nilpotent) normal subgroup of $G$. Both these subgroup are characteristic in $G$. Recall that $G$ is said to be semisimple if its radical is trivial. In particular, $G/R$ is semisimple unless $G$ is solvable. For a closed subgroup $H$ of $G$, let $H^{0}$ denote the connected component of the identity $e$ in $H$, $[H,H]$ denote the commutator subgroup of $H$. Note that $H^{0}$, $[H,H]$ and its closure $\overline{[H,H]}$ are characteristic in $H$. Let ${\mathfrak{g}}$ denote the Lie algebra of the connected Lie group $G$. There is an exponential map $\exp:{\mathfrak{g}}\to G$ which is a continuous map and its restriction to a small open neighbourhood of 00 in ${\mathfrak{g}}$ is a homeomorphism onto an open neighbourhood of the identity $e$ in $G$. Given $T\in{\rm Aut}(G),$ there exists a corresponding Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp\circ\,{\rm d\,}T=T\circ\,\exp$. The map ${\rm d\,}:{\rm Aut}(G)\rightarrow{\rm Aut}({\mathfrak{g}})$, $T\mapsto{\rm d\,}T$ is an injective continuous homomorphism and its image is closed, and in fact, ${\rm d\,}$ is a homeomorphism onto its image (cf. [9], see also [22]) It is known that ${\rm Aut}({\mathfrak{g}})$ is a closed subgroup of ${\rm GL}({\mathfrak{g}})$, and we can view ${\rm Aut}(G)$ as a closed subgroup of ${\rm GL}({\mathfrak{g}})$. The topology on ${\rm Aut}(G)$, when considered as a subspace topology of ${\rm GL}({\mathfrak{g}})$, coincides with the (modified) compact-open topology (see [44]). We refer the reader to [23] for the basic structure theory of Lie groups.
In §2, we study general properties of ${\rm Sub}_{G}$ and prove some elementary results about the structure of ${\rm Sub}^{p}_{G}$. In §3, we discuss some properties and actions of automorphisms of a connected Lie group $G$ on $G$ as well as ${\rm Sub}^{p}_{G}$ and discuss some preliminary results. In §4 we prove the main results stated above.
2. Properties and structure of ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some basic properties of the Chabauty topology on the space ${\rm Sub}_{G}$. We also discuss some elementary results about the structure of ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$.
Let $L$ be a locally compact Hausdorff topological group. For a compact set $K$ and an open set $U$ in $L$, let $U_{1}(K)=\{H\in{\rm Sub}_{L}\mid H\cap K=\emptyset\}$ and $U_{2}(U)=\{H\in{\rm Sub}_{L}\mid H\cap U\neq\emptyset\}$. Here, $\{U_{1}(K)\mid K\mbox{ is compact}\}\cup\,\{U_{2}(U)\mid U\mbox{ is open}\}$ is a sub-basis of the Chabauty topology on ${\rm Sub}_{L}$. As mentioned earlier, ${\rm Sub}_{L}$ is compact and Hausdorff, and if $L$ is second countable, then ${\rm Sub}_{L}$ is metrisable (cf.  [6], Lemma E.1.1). In particular ${\rm Sub}_{L}$ is metrisable if $L$ is a connected Lie group. For a closed subgroup $H$ of $L$, it is easy to see that ${\rm Sub}_{H}$ carries the subspace topology of ${\rm Sub}_{L}$. The following lemma is known (see e.g. Proposition E.1.2 in [6])
Lemma 2.1.
Let $G$ be a connected Lie group. A sequence $\{H_{n}\}\subset{\rm Sub}_{G}$ converges to $H\in{\rm Sub}_{G}$ if and only if the following hold:
  1. ${\rm(I)}$
    For $g\in G$, if there exists a subsequence $\{H_{n_{k}}\}$ of $\{H_{n}\}$ with $h_{k}\in H_{n_{k}}$, $k\in{\mathbb{N}}$, such that $h_{k}\to g$ in $G$, then $g\in H$.
  2. ${\rm(II)}$
    For every $h\in H$, there exists a sequence $\{h_{n}\}_{n\in{\mathbb{N}}}$ such that $h_{n}\in H_{n}$, $n\in{\mathbb{N}}$, and $h_{n}\to h$.
The following lemma is a consequence of the well-known result that the set of discrete subgroups of a connected Lie group $G$ is open in ${\rm Sub}_{G}$. We will give a short proof for the sake of completeness.
Lemma 2.2.
Let $G$ be a connected Lie group. Then the trivial subgroup $\{e\}$ is isolated in ${\rm Sub}^{co}_{G}$, and so in ${\rm Sub}^{p}_{G}$. Moreover, ${\rm Sub}^{co}_{G}$ does not contain any non-trivial discrete subgroup.
Proof.
Let $\mathfrak{C}$ be the set of closed connected subgroups of $G$. We have ¯$\overline{\mathfrak{C}}={\rm Sub}^{co}_{G}$. By Proposition E.1.5 in [6] (see also Proposition 3.4(iii) in [8], or also Corollary of Lemma 1.2 in [45]) the set $\mathfrak{D}$ of discrete subgroups of $G$ is open in ${\rm Sub}_{G}$ and so $\{\{e\}\}\subseteq\mathfrak{D}\cap{\rm Sub}^{co}_{G}=\mathfrak{D}\cap\overline{\mathfrak{C}}\subseteq\overline{\mathfrak{D}\cap\mathfrak{C}}=\{\{e\}\}$. Hence the assertions hold. ∎
Note that the action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ is continuous for connected Lie groups $G$ (cf. [42], Lemma 2.4). In particular, if $T$ generates a relatively compact group in ${\rm Aut}(G)$, then it acts distally on $G$ as well as on ${\rm Sub}_{G}$ (see e.g. [31], Lemma 4.2 for the latter statement).
A (continuous real) one-parameter subgroup in $G$ is the image of a continuous homomorphism from $\mathbb{R}$ to $G$. Any one-parameter subgroup is of the form $\{\exp tX\}_{t\in{\mathbb{R}}}$ for some $X$ in the Lie algebra ${\mathfrak{g}}$ of $G$. Moreover, if $H$ is a one-parameter subgroup in $G$, then so is $T(H)$ for any $T\in{\rm Aut}(G)$. Let
${\mathcal{P}}_{1}(G):=\{A\mid A\mbox{ is a closed one-parameter subgroup of }G\}.$
Now ${\rm Sub}^{p}_{G}$ is the closure (in ${\rm Sub}_{G}$) of the set ${\mathcal{P}}_{1}(G)$, i.e. ${\rm Sub}^{p}_{G}:=\overline{{\mathcal{P}}_{1}(G)}$, which is compact and invariant under the action of ${\rm Aut}(G)$.
Note that if any one-parameter subgroup in a connected Lie group is not closed, then its closure is compact and isomorphic to an $n$-dimensional torus $\mathbb{T}^{n}=(\mathbb{S}^{1})^{n}$ for some $n\in{\mathbb{N}}$ with $n\geq 2$. If it is closed and bounded (compact), then it is isomorphic to $\mathbb{S}^{1}$. If it is unbounded (not relatively compact), then it is closed and isomorphic to ${\mathbb{R}}$.
For $\{x_{t}\}_{t\in{\mathbb{R}}}$, which may be a continuous one-parameter set or a one-parameter subgroup, we may just write $\{x_{t}\}$ for brevity, i.e. we may omit the parameter set if there is no ambiguity.
Now we prove some elementary results about the structure of ${\rm Sub}^{p}_{G}$. The following lemma shows that if $G$ contains any non-trivial compact connected abelian subgroup of dimension at least 2, then ${\mathcal{P}}_{1}(G)$ is not closed in ${\rm Sub}_{G}$.
Lemma 2.3.
Let $G$ be a connected Lie group. Then the following hold:
  1. (1)
    All compact connected abelian subgroups of $G$ belong to ${\rm Sub}^{p}_{G}$.
  2. (2)
    If $\{x_{t}\}\in{\mathcal{P}}_{1}(G)\subset{\rm Sub}^{p}_{G}$ and if $K$ is a compact connected abelian subgroup of $G$ which is centralised by each $x_{t}$. Then $\{x_{t}\}K\in{\rm Sub}^{p}_{G}$.
Proof.
Let $K$ be a compact connected abelian subgroup of $G$. Then the set $F$ of all finite order elements in $K$ is dense in $K$ and each such element of $F$ is contained in a closed one-parameter subgroup of $K$. Since $K$ is monothetic (i.e. it has a dense cyclic subgroup), we get that $K$ is a limit of a sequence of compact one-parameter subgroups of $K$ in ${\rm Sub}_{K}$. Therefore, $K\in{\rm Sub}^{p}_{K}\subset{\rm Sub}^{p}_{G}$. Thus (1) holds.
Let $\{x_{t}\}\in{\mathcal{P}}_{1}(G)\subset{\rm Sub}^{p}_{G}$ be such that each $x_{t}$ centralises $K$. If $\{x_{t}\}$ is compact, then $\{x_{t}\}K$ is a compact connected abelian subgroup of $G$ and hence $\{x_{t}\}K\in{\rm Sub}^{p}_{G}$ from (1). Now suppose $\{x_{t}\}$ is isomorphic to ${\mathbb{R}}$. Let $k\in K$ be such that it generates a dense cyclic subgroup in $K$. Since $K$ is exponential, there exists a one-parameter subgroup $\{k_{t}\}$, which is dense in $K$ and $k=k_{1}$. Let $A_{n}=\{a_{t}(n)=x_{t/n}k_{t}\}_{t\in{\mathbb{R}}}$. Then each $A_{n}$ is a closed one-parameter subgroup in $G$, and passing to a subsequence if necessary, we get that $A_{n}\to A$ for some subgroup $A\in{\rm Sub}^{p}_{G}$. Note that $\{x_{t}\}\times K$ is a closed subgroup of $G$ and it contains each $A_{n}$. Therefore, $A\subset\{x_{t}\}\times K$. Let $t\in{\mathbb{R}}$ be fixed. Since $a_{t}(n)=x_{t/n}k_{t}\to k_{t}$ as $n\to\infty$, we get that $k_{t}\in A$, and hence $K\subset A$. Let $b_{n}=a_{tn}(n)=x_{t}k_{tn}\in A_{n}$, $n\in{\mathbb{N}}$. Passing to a further subsequence of $\{b_{n}\}$ if necessary, we get that $b_{n}\to x_{t}h$ for some $h\in K$, and hence, $x_{t}\in A$, as $K\subset A$. Now $\{x_{t}\}\subset A$. Hence, $A=\{x_{t}\}\times K$ and (2) holds. ∎
The following lemma relates the set ${\rm Sub}^{p}_{G/K}$ and a compact subset of ${\rm Sub}^{p}_{G}$ which is invariant under the action of ${\rm Aut}(G)$.
Lemma 2.4.
Let $G$ be a connected Lie group, $K$ be a compact connected central subgroup of $G$ and let $\pi:G\to G/K$ be the natural projection. Then the following hold:
  1. (1)
    $\{HK\mid H\in{\rm Sub}_{G}\}$ is a closed (compact) subset of ${\rm Sub}_{G}$ and $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$ is a closed (compact) subset of ${\rm Sub}^{p}_{G}$, and they are $T$-invariant if $T(K)=K$, where $T\in{\rm Aut}(G)$.
  2. (2)
    ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$.
  3. (3)
    ${\rm Sub}^{p}_{G/K}=\{\pi(H)\mid H\in{\rm Sub}^{p}_{G}\}$.
  4. (4)
    ${\rm Sub}^{p}_{G/K}$ is homeomorphic to $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$, and the latter is the same set as $\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$.
Proof.
Let $K$ and $\pi$ be as above. Note that $K\in{\rm Sub}^{p}_{G}$ by Lemma 2.3. If $G=K$, then it is easy to see that $(1-4)$ are trivially satisfied.
Now suppose that $G\neq K$. Since $K$ is compact and central in $G$, $HK$ is a closed subgroup of $G$ for any closed subgroup $H$ of $G$. If $H_{n}K\to L$ in ${\rm Sub}_{G}$, then since both ${\rm Sub}_{G}$ and $K$ are compact, it is easy to show that for any limit point $H$ of $\{H_{n}\}$, $L=HK$. Therefore, $\{HK\mid H\in{\rm Sub}_{G}\}$ is closed. Now if $H\in{\rm Sub}^{p}_{G}$, then there exists $\{A_{n}\}\subset{\mathcal{P}}_{1}(G)$, such that $A_{n}\to H$. By Lemma 2.3 (2), $A_{n}K\in{\rm Sub}^{p}_{G}$. As $A_{n}K\to HK$ and ${\rm Sub}^{p}_{G}$ is closed, we get that $HK\in{\rm Sub}^{p}_{G}$. Now arguing as above, it is easy to show that $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$ is a closed subset of ${\rm Sub}^{p}_{G}$. Moreover, if $T\in{\rm Aut}(G)$ with $T(K)=K$, then $T(HK)=T(H)K$ for any $H\in{\rm Sub}_{G}$. Therefore, the sets mentioned in (1) are invariant under the action of ${\rm Aut}(G)$ if $T(K)=K$. Thus (1) holds.
If $A\in{\mathcal{P}}_{1}(G)$, then $AK$ is closed and $\pi(A)\in{\mathcal{P}}_{1}(G/K)$. Conversely, suppose $B\in{\mathcal{P}}_{1}(G/K$) is nontrivial. It is easy to see that $B_{1}:=\pi^{-1}(B)$ is a closed connected nilpotent Lie subgroup of $G$. There exists a neighbourhood $U$ of $e$ such that every element of $U$ is contained in a one-parameter subgroup and $U\setminus K$ is nonempty. Choose any $x\in U\setminus K$, there exists a one-parameter subgroup $\{x_{t}\}$ with $x_{1}=x$. Then $\pi(\{x_{t}\})=B$ and $B_{1}=\{x_{t}\}K$. As $K$ is central in $G$ we get that $B_{1}$ is, in fact, abelian. If $\{x_{t}\}$ is closed, then it belongs to ${\mathcal{P}}_{1}(G)$ and $\pi(\{x_{t}\})=B$. Now suppose it is not closed, then its closure is compact, and hence $B_{1}$ is compact. This implies that the set of torsion elements is dense in $B_{1}$. Since $U\setminus K$ is open, we can choose $y\in B_{1}\cap(U\setminus K)$ such that $y^{n}=e$, for some $n\in{\mathbb{N}}\setminus\{1\}$. Let $\{y_{t}\}$ be a one-parameter subgroup in $B_{1}$ such that $y_{1}=y$. Then $\pi(\{y_{t}\})=B$. Here, $\{y_{t}\}$ is closed as $y_{1}^{n}=e$. Therefore, ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$, i.e. (2) holds.
Let $\pi^{\prime}:{\rm Sub}^{p}_{G}\to{\rm Sub}^{p}_{G/K}$ be the map induced by $\pi$. Then $\pi^{\prime}$ is a continuous closed map. From (2), we get that ${\rm Sub}^{p}_{G/K}=\overline{{\mathcal{P}}_{1}(G/K)}=\pi^{\prime}(\overline{{\mathcal{P}}_{1}(G)})=\pi^{\prime}({\rm Sub}^{p}_{G})=\{\pi(H)\mid H\in{\rm Sub}^{p}_{G}\}$. Thus (3) holds.
Let ${\mathcal{S}}(K):=\{HK\mid H\in{\rm Sub}^{p}_{G}\}$. By (1), ${\mathcal{S}}(K)$ is closed in ${\rm Sub}^{p}_{G}$, and hence it is compact. Observe that ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$. By (3), $\pi^{\prime}({\mathcal{S}}(K))=\pi^{\prime}({\rm Sub}^{p}_{G})={\rm Sub}^{p}_{G/K}$ and it is obvious that $\pi^{\prime}|_{{\mathcal{S}}(K)}$ is continuous and bijective, and hence it is a homeomorphism. Thus (4) holds. ∎
3. Some properties of automorphisms and their actions on $G$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some properties and actions of automorphisms of a connected Lie group $G$ on $G$ as well as ${\rm Sub}^{p}_{G}$ and discuss some preliminary results which are useful.
For convenience, we introduce a class ${\mathcal{C}^{\prime}}$ of Lie groups as follows:
Definition 3.1.
A connected Lie group $G$ is said to be in class ${\mathcal{C}^{\prime}}$ if $G$ has no compact central subgroup of positive dimension ((((i.e. if $G$ has no non-trivial compact connected central subgroup)))).
The notation is inspired by the fact that the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$ defined by Dani and McCrudden in [13]. The class $\mathcal{C}$ consists of connected Lie groups which admit a (real) linear representation with discrete kernel. Proposition 2.5 of [13] shows that $G$ is in class $\mathcal{C}$ if and only if for the radical $R$ of $G$, the intersection of $\overline{[R,R]}$ with the center of $G$ has no non-trivial compact subgroup. Therefore, the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$. Note that any connected Lie group $G$ admits a maximal compact connected central subgroup $M$, and $G/M$ has no non-trivial compact connected central subgroup and hence it belongs to class ${\mathcal{C}^{\prime}}$. We may recall a result of Dani in [12] that if $G$ belongs to class ${\mathcal{C}^{\prime}}$, then ${\rm Aut}(G)$ is almost algebraic, i.e. it is a(n open) subgroup of finite index in an algebraic subgroup of ${\rm GL}({\mathfrak{g}})$, and in particular, it has finitely many connected components.
It is known that $T\in{\rm Aut}(G)$ acts distally on $G$ if and only if ${\rm d\,}T$ acts distally on the Lie algebra ${\mathfrak{g}}$ of $G$, and the latter statement is equivalent to the statement that all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf. [2]). It is also known that if $G$ admits an automorphism $T$ such that ${\rm d\,}T$ has no eigenvalue of absolute value 1 (equivalently, if $T$ is expansive; see [38], and also [7]), then $G$ must be nilpotent (this is well-known, see e.g. [38]). We will need the following lemma which is slightly stronger. The lemma may be known, but we will give a short proof for the sake of completeness.
Lemma 3.2.
Let $G$ be a connected non-nilpotent Lie group and let $T$ be an automorphism of $G$. Then at least one eigenvalue of ${\rm d\,}T$ is a root of unity.
Proof.
Let $M$ be the largest compact connected central subgroup of $G$. Then $G/M$ is in class ${\mathcal{C}^{\prime}}$. As $G$ is not nilpotent, $G/M$ is also not nilpotent. Since $M$ is characteristic in $G$, we may replace $G$ by $G/M$ and assume that $G\in{\mathcal{C}^{\prime}}$. If $T$ has finite order, then $({\rm d\,}T)^{n}={\rm Id}$ for some $n\in{\mathbb{N}}$, and the assertion follows in this case. Now suppose $T$ does not have finite order.
Since $G\in{\mathcal{C}^{\prime}}$, ${\rm Aut}(G)$ is almost algebraic (cf. [12]), and hence it has finitely many connected components. Now we may replace $T$ by its suitable power and assume that $T\in{\rm Aut}(G)^{0}$, the connected component of the identity in ${\rm Aut}(G)$, and that $T$ is contained in a non-trivial one-parameter subgroup (say) $\{T_{t}\}$ of ${\rm Aut}(G)^{0}$ as $T=T_{1}$. We will show that 1 is an eigenvalue of ${\rm d\,}T$.
Suppose $G$ is solvable. Let $H=\{T_{t}\}\ltimes G$. Then $H$ is solvable and the nilradical of $H$, being contained in $G$, is the same as the nilradical $N$ of $G$. Therefore, $H/N$ is abelian. Thus $T$ acts trivially on $G/N$. As $G$ is not nilpotent, $G\neq N$, and hence 1 is an eigenvalue ${\rm d\,}T$.
Now suppose $G$ is not solvable. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$, and let $\overline{T}_{t}$ be the automorphisms of $G/R$ induced by $T_{t}$ for each $t$. It is enough to prove that 1 is an eigenvalue of ${\rm d\,}\overline{T}$. Hence we may replace $G$ by $G/R$ and $\{T_{t}\}$ by $\{\overline{T}_{t}\}$ and assume that $G$ is semisimple. Now ${\rm Aut}(G)^{0}$ consists of inner automorphisms of $G$ and $T_{t}={\rm inn}(x_{t})$, $t\in{\mathbb{R}}$, for some one-parameter subgroup $\{x_{t}\}\subset G$. If $x_{1}=e$, then $T=T_{1}={\rm Id}$. Now suppose $x_{1}\neq e$. Then $\{x_{t}\}$ is non-trivial. Let $X\in{\mathfrak{g}}$ be such that $\exp tX=x_{t}$, $t\in{\mathbb{R}}$. As $T(x_{t})=x_{t}$, $t\in{\mathbb{R}}$, we have that ${\rm d\,}T(X)=X$ and 1 is an eigenvalue of ${\rm d\,}T$. ∎
Remark 3.3.
The proof above also shows that for a connected solvable Lie group $G$ with the nilradical $N$, for any $T\in{\rm Aut}(G)$, $\overline{T}$ has finite order, where $\overline{T}$ is the automorphism of $G/N$ induced by $T$. In particular, all the eigenvalues of ${\rm d\,}\overline{T}$ on the Lie algebra of $G/N$ are roots of unity.
For a connected Lie group $G$ and $T\in{\rm Aut}(G)$, let $C(T)$ denote the contraction group of $T$ and it is defined as
$C(T):=\{x\in G\mid T^{n}(x)\to e\mbox{ as }n\to\infty\}.$
Note that $C(T)$ is a connected nilpotent subgroup of $G,$ and if it is closed in $G$, then it is simply connected (see [43], see also [21]). Both $C(T)$ and $C(T^{-1})$ are trivial if and only if $T$ acts distally on $G$ (cf.  [35], Theorem 4.1). Note also that $C(T)$ is trivial if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value greater than or equal to 1 (cf. [35, 2]).
We will need the following lemma which can be proven easily by using Theorem 1.1 of [14], Theorem 2.4 and Corollary 2.7 of [21]; we will give a sketch of proof. Note that the largest compact connected central subgroup of $G$ is characteristic in $G$. The lemma generalises a part of the statement in Proposition 4.3 of [35].
Lemma 3.4.
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. Then the following are equivalent:
  1. (1)
    $T$ acts distally on the largest compact connected central subgroup of $G$.
  2. (2)
    The contraction groups $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent subgroups of $G$.
  3. (3)
    $T$ acts distally on every compact $T$-invariant subgroup of $G$.
Proof.
Let $M$ be the largest compact connected central subgroup of $G$. If $G=M$, then it is easy to see that $(1-3)$ are equivalent. Now suppose $G\neq M$. Let $\pi:G\to G/M$ be the natural projection and let $T^{\prime}\in{\rm Aut}(G/M)$ be the automorphism induced by $T$ on $G/M$. As $G/M\in{\mathcal{C}^{\prime}}$, using Theorem 1.1 of [14], one can show that $C(T^{\prime})$ is closed (see the proof of Proposition 4.3 of [35] for more details). Hence $\pi^{-1}(C(T^{\prime}))=C_{M}(T):=\{x\in G\mid T^{n}(x)M\to M\mbox{ in }G/M\}$ is closed. Suppose (1) holds, i.e. $T$ acts distally on $M$. Then $C(T)\cap M=C(T|_{M})=\{e\}$. As $C_{M}(T)$ is closed, by Corollary 2.7 of [21], $C(T)$ is closed. Similarly, $C(T^{-1})$ is closed. The rest of the assertion in (2) follows from Corollary 2.4 in [43]. This proves $(1)\implies(2)$.
Now suppose (2) holds, i.e. $C(T)$ is closed. Let $K$ be any compact $T$-invariant subgroup of $G$. Then $C(T|_{K})=C(T)\cap K$ is a compact subgroup of $C(T)$. Hence $C(T|_{K})$ is trivial (cf. [21], Corollary 2.5). Similarly, $C((T|_{K})^{-1})$ is trivial and $T$ acts distally on $K$ (cf. [26]). Thus (3) holds. $(3)\implies(1)$ is obvious. ∎
We note the following useful proposition which will enable us to work on a quotient Lie group which belongs to class ${\mathcal{C}^{\prime}}$.
Proposition 3.5.
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. Let $K$ be a $T$-invariant compact connected central subgroup of $G$ and let $\overline{T}$ denote the automorphism of $G/K$ induced by $T$. If $T$ acts distally on ${\rm Sub}^{p}_{G}$, then $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/K}$.
Proof.
Let $T\in{\rm Aut}(G)$ and let $\pi:G\to G/K$ be the natural projection. Here, $\overline{T}(\pi(H))=\pi(T(H))$ for any $H\in{\rm Sub}_{G}$, and ${\rm Sub}^{p}_{G/K}$ is $\overline{T}$ invariant. Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=H\}$. It is a compact $T$-invariant subset of ${\rm Sub}^{p}_{G}$ and it is homeomorphic to ${\rm Sub}^{p}_{G/K}$ under the map induced by $\pi$ on ${\mathcal{S}}(K)$ (cf. Lemma 2.4 (4)). Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then the $T$-action on ${\mathcal{S}}(K)$ is distal. As $\pi\circ T=\overline{T}\circ\pi$, we get that $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/K}$. ∎
We will also need the following elementary lemma.
Lemma 3.6.
Let $G$ be a simply connected nilpotent Lie group with the Lie algebra ${\mathfrak{g}}$ and let $T\in{\rm Aut}(G)$. Then the set ${\mathcal{P}}_{1}(G)$ consisting of all closed one-parameter subgroups of $G$ is closed in ${\rm Sub}_{G}$ and is the same as ${\rm Sub}^{p}_{G}$. Moreover, $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$.
Proof.
In a simply connected nilpotent Lie group, all one-parameter subgroups are closed and there is a one-to-one correspondence between non-trivial (closed) one-parameter subgroups of $G$ and one-dimensional subspaces of ${\mathfrak{g}}$. The correspondence is via the exponential map $\exp:{\mathfrak{g}}\to G$ with $\log$ as its inverse. Namely; for $\{x_{t}\}$, $x_{t}=\exp tX$, for some $X\in{\mathfrak{g}}$, and $\log x_{t}=tX$, $t\in{\mathbb{R}}$. Also the exponential map induces a homeomorphism from ${\mathcal{P}}_{1}({\mathfrak{g}})\to{\mathcal{P}}_{1}(G)$. As $\exp\circ\,{\rm d\,}T=T\circ\exp$, it follows that $T$ acts distally on ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$ if and only if ${\rm d\,}T$ acts distally on ${\mathcal{P}}_{1}({\mathfrak{g}})={\rm Sub}^{p}_{\mathfrak{g}}$. ∎
4. Proofs of the main results
In this section, we prove the main results stated in the introduction. Additionally, we get some corollaries which are derived from the main results. We give some examples of Lie groups $G$ with a nontrivial central torus on which all automorphisms act distally, which is stronger than one of the conditions in Theorems 1.2, 1.4 and 1.5.
Proof of Theorem 1.1 $(1)$ and $(2)$ for nilpotent groups.
Let $G$ be a (non-trivial) connected Lie group without any non-trivial compact connected central subgroup, i.e. $G$ belongs to class ${\mathcal{C}^{\prime}}$. Let $T\in{\rm Aut}(G)$.
[…]
Conversely, for $G={\mathbb{R}}^{d}$ and $T\in{\mathcal{K}}{\mathcal{D}}$, with a compact subgroup ${\mathcal{K}}$ of ${\rm GL}(d,{\mathbb{R}})$ and the center ${\mathcal{D}}$ of ${\rm GL}(d,{\mathbb{R}})$, the action of $T$ on ${\rm Sub}^{p}_{G}$ is same as that of $S$ on ${\rm Sub}^{p}_{G}$ for some $S\in{\mathcal{K}}$ since ${\mathcal{D}}$ acts trivially on ${\rm Sub}^{p}_{G}$. As the compact subgroups of ${\rm Aut}(G)$ act distally on any closed invariant subspace of ${\rm Sub}_{G}$ (${\rm Sub}^{p}_{G}$ in this case), we have that $T$ acts distally on ${\rm Sub}^{p}_{G}$. This completes the proof of Theorem 1.1 (1).
$(2):$ Suppose $G$ belongs to the class ${\mathcal{C}^{\prime}}$ and it is non-abelian and nilpotent. Then $G$ is simply connected as $G$ has no compact central subgroup of positive dimension. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. We want to show that $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. Since $G$ belongs to the class ${\mathcal{C}^{\prime}}$, ${\rm Aut}(G)$ is almost algebraic as a subgroup of ${\rm GL}({\mathfrak{g}})$ (cf. [12]).
As $G$ is simply connected and nilpotent, by Lemma 3.6, ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$. Then from (1) above, we have that ${\rm d\,}T\in{\mathcal{K}}\{r\,{\rm Id}\mid r\neq 0\}$ for some compact subgroup ${\mathcal{K}}$ of ${\rm GL}({\mathfrak{g}})$ and ${\rm Id}$ is the identity map on ${\mathfrak{g}}$. Here, ${\rm d\,}T=\psi\,\circ\,r\,{\rm Id}$ for some $r\in{\mathbb{R}}\setminus\{0\}$ and $\psi\in{\mathcal{K}}$. If $r=\pm 1$, then ${\rm d\,}T^{2}\in{\mathcal{K}}$ and $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. If possible suppose $r\neq\pm 1$. Since ${\rm Aut}(G)$ is almost algebraic, there exists $n\in{\mathbb{N}}$ such that ${\rm d\,}T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathbb{R}}\times\mathbb{T}^{k}$ of ${\rm Aut}(G)^{0}$ (which is identified with its image in ${\rm GL}({\mathfrak{g}})$), where $\mathbb{T}^{k}=(\mathbb{S}^{1})^{k}$ for some $k\in{\mathbb{N}}$. Therefore, $\psi^{n}\,\circ\,r^{n}\,{\rm Id}\in{\mathbb{R}}\times\mathbb{T}^{k}\subset{\rm Aut}(G)^{0}$. Hence both $\psi^{n}$ and $r^{n}\,{\rm Id}\in{\rm Aut}(G)^{0}$. But $r^{n}\,{\rm Id}$ is not a Lie algebra automorphism of ${\mathfrak{g}}$ as $r^{n}\neq 1$; otherwise $r^{n}[X,Y]=r^{n}{\rm Id}[X,Y]=[r^{n}{\rm Id}(X),r^{n}{\rm Id}(Y)]=r^{2n}[X,Y]$, and it would imply that $[X,Y]=0$ for all $X,Y\in{\mathfrak{g}}$, which leads to a contradiction as $G$ is not abelian. Thus, $r=\pm 1$ and $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse statement is obvious. Thus (2) holds if $G$ is nilpotent. ∎
We will continue with the proof of Theorem 1.1 after proving Theorems 1.2 and 1.3.
Proof of Theorem 1.2.
Let $G$ be a connected Lie group which is not a vector group. Let $M$ be the largest compact connected central subgroup of $G$. Suppose $T\in{\rm Aut}(G)$ be such that it acts distally on $M$. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ acts distally on $G$. The assertion holds trivially if $G=M$.
Step 1: Suppose $G$ is simply connected and nilpotent. Since $G$ is not a vector group, $G$ is non-abelian and by the proof of Theorem 1.1 (2) for the nilpotent case, $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. Hence $T$ acts distally on $G$.
Step 2: Now suppose $G$ is not a simply connected nilpotent group. By Theorem 4.1 of [35], it is enough to prove that the contraction groups $C(T)$ and $C(T^{-1})$ are trivial. As $T$ acts distally on the maximal compact connected central subgroup $M$, by Lemma 3.4, both $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent subgroups. Suppose $C(T)$ is non-trivial. If possible, suppose $C(T)$ is non-abelian. As it is $T$-invariant and $T|_{C(T)}$ acts distally on ${\rm Sub}^{p}_{C(T)}$, we get from Step 1 that $T|_{C(T)}$ acts distally on $C(T)$, a contradiction. Thus $C(T)$ is abelian and $C(T)$ is isomorphic to ${\mathbb{R}}^{d}$ as it is simply connected. By Theorem 1.1 (1), $T_{1}=T|_{C(T)}=\phi\circ\alpha\,I$ for some $\phi$ contained in a compact subgroup of ${\rm GL}(d,{\mathbb{R}})$ and $\alpha\in{\mathbb{R}}$, $10<|\alpha|<$ 1; here $I$ denotes the identity matrix in ${\rm GL}(d,{\mathbb{R}})$, replacing $T$ by $T^{2}$, we may assume that $\alpha>0$.
Step 3: Suppose $G$ is nilpotent but not simply connected. Then $M$ as above is non-trivial. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], we have that $T^{k}|_{M}$ is unipotent for some $k\in{\mathbb{N}}$, i.e. ${\rm d\,}T$ has an eigenvalue which is a root of unity. If $G$ is not nilpotent, by Lemma 3.2, ${\rm d\,}T$ has an eigenvalue which is a root of unity. We may replace $T$ by $T^{m}$, for some $m\in{\mathbb{N}}$, and assume that 1 is an eigenvalue of ${\rm d\,}T$. Let $X\in{\mathfrak{g}}$, $X\neq 0$ be such that ${\rm d\,}T(X)=X$ and let $A=\{\exp tX\}$, a one-parameter subgroup, it is either closed or its closure $\bar{A}$ is compact. Since $T|_{A}={\rm Id}$, we have that $T|_{\bar{A}}={\rm Id}$ and it follows that $\bar{A}$ normalises $C(T)$ and $\bar{A}\cap C(T)=\{e\}$.
We now show that $C(T)$ normalises $\bar{A}$, which would imply that $\bar{A}\,C(T)$ is abelian as $C(T)$ and $A$ are so. Here, $T(\bar{A})=\bar{A}$ and $\bar{A}\in{\rm Sub}^{p}_{G}$. Now for any nonzero $c\in C(T)$, we have that
$T^{n}(c\bar{A}c^{-1})=T^{n}(c)\,\bar{A}\,T^{n}(c^{-1})\to\bar{A}.$
Since $T$ acts distally on ${\rm Sub}^{p}_{G}$, we get that $c\bar{A}c^{-1}=\bar{A}$, for all $c\in C(T)$. Thus $C(T)$ normalises $\bar{A}$, and hence $\bar{A}\,C(T)$ is abelian.
Let $H=\overline{A\times C(T)}$. Then $H$ is a closed $T$-invariant abelian subgroup and $\bar{A}\subset H$. First, suppose $H$ is a vector group. By Theorem 1.1 (1), $T|_{H}\in{\mathcal{K}}{\mathcal{D}}$, for some compact subgroup ${\mathcal{K}}$ in ${\rm GL}(H)$ and the center ${\mathcal{D}}$ of ${\rm GL}(H)$, and thus all the eigenvalues of $T|_{H}$ have the same absolute value. Since $T|_{\bar{A}}={\rm Id}$ and $T|_{C(T)}$ have eigenvalues of absolute value $\alpha\neq 1$, it leads to a contradiction. Thus $C(T)$ must be trivial in this case.
Step 4: Now suppose $H$ is not a vector group. Since $C(T)$ is closed and the image of $A$ is dense in $H/C(T)$, we get that $H=K\times C(T)$ for the maximal compact subgroup $K$ of $H$. As $T$ acts trivially on $H/C(T)$ and $K$ is $T$-invariant, we get that $T|_{K}={\rm Id}$.
Let $\psi,S:H\to H$ be such that $S|_{K}={\rm Id}$, $S|_{C(T)}=\alpha\,{\rm Id}$, $\psi|_{K}={\rm Id}$ and $\psi|_{C(T)}=\phi$, where $\alpha$ and $\phi$ are as in Step 2. Now $T|_{H}=S\psi=\psi S$, and $\psi$ is contained in a compact subgroup of ${\rm Aut}(H)$. Therefore, $S$ acts distally on ${\rm Sub}^{p}_{H}$ (cf. [41], Lemma 2.2), and $C(S)=C(T)$.
Let $\{z_{t}=x_{t}c_{t}\}$ be a closed one-parameter subgroup, where $\{x_{t}\}\subset K$ is closed and isomorphic to ${\mathbb{S}}^{1}$ and $\{c_{t}\}\subset C(S)=C(T)$ is a non-trivial (closed) one-parameter subgroup such that $S(c_{t})=c_{\alpha t}$, $t\in{\mathbb{R}}$ (as addition is the group operation on $C(T)={\mathbb{R}}^{d}$, we choose $c_{t}=tc$, $t\in{\mathbb{R}}$, for some nonzero $c\in{\mathbb{R}}^{d}$). Let $\{n_{k}\}$ be an unbounded sequence in ${\mathbb{N}}$ such that $S^{n_{k}}(\{z_{t}\})\to C$ for some $C$ in ${\rm Sub}^{p}_{H}$. Note that $\{z_{t}\}\subset\{x_{t}\}\times\{c_{t}\}$ and the latter is closed and $S$-invariant. So $C\subset\{x_{t}\}\times\{c_{t}\}$. For a fixed $t\in{\mathbb{R}}$, as $c_{t}\in C(S)=C(T)$, we get that
$S^{n_{k}}(z_{t})=x_{t}S^{n_{k}}(c_{t})\to x_{t},$
as $n_{k}\to\infty$. Therefore, $\{x_{t}\}\subset C$. Again for a fixed $t\in{\mathbb{R}}$,
$S^{n_{k}}(z_{\alpha^{-n_{k}}t})=x_{\alpha^{-n_{k}}t}S^{n_{k}}(c_{\alpha^{-n_{k}}t})=x_{\alpha^{-n_{k}}t}c_{t}\to x_{s}c_{t}$
along a subsequence of $\{n_{k}\}$, for some $s\in{\mathbb{R}}$. Therefore, $x_{s}c_{t}$, and hence, $c_{t}$ belongs to $C$. Thus $\{c_{t}\}\subset C$ and $C=\{x_{t}\}\times\{c_{t}\}$, and hence $C$ is $S$-invariant. This implies that $S^{n_{k}}(S(\{z_{t}\}))\to C$, and hence that $S(\{z_{t}\})=\{z_{t}\}$ since $S$ acts distally on ${\rm Sub}^{p}_{G}$. As $\{z_{t}\}$ is isomorphic to ${\mathbb{R}}$, there exists $\beta\in{\mathbb{R}}\setminus\{0\}$ such that for all $t\in{\mathbb{R}}$,
$S(z_{t})=z_{\beta t}=x_{\beta t}c_{\beta t}\ \mbox{ and }\ S(z_{t})=S(x_{t}c_{t})=x_{t}c_{\alpha t}.$
As $\{x_{t}\}\times\{c_{t}\}$ is a direct product and $\{c_{t}\}$ is isomorphic to ${\mathbb{R}}$, it follows that $\beta=\alpha$ and $x_{t}=x_{\alpha t}$, i.e. $x_{(1-\alpha)t}=e$ for all $t\in{\mathbb{R}}$. As $\alpha\neq 1$, $x_{t}=e$ for each $t$, a contradiction. Therefore, $C(T)$ is trivial.
Similarly, we can show that $C(T^{-1})$ is trivial, and hence $T$ acts distally on $G$. ∎
In view of Theorem 1.1 of [2], the following corollary is an easy consequence of Theorem 1.2. We omit the proof.
Corollary 4.1.
Let $G$ be a connected Lie group which is not a vector group and let $H$ be a subgroup of ${\rm Aut}(G)$ such that it acts distally on the largest compact connected central subgroup of $G$. If $H$ acts distally on ${\rm Sub}^{p}_{G}$, then $H$ acts distally on $G$.
Before proving Theorem 1.3, we first prove the following proposition which is a special case of the theorem.
Proposition 4.2.
Let $G$ be a connected Lie group, $T\in{\rm Aut}(G)$ and let $L$ be a closed connected nilpotent normal $T$-invariant subgroup of $G$. Suppose $T$ acts trivially on both $L$ and $G/L$. If $T$ acts distally on ${\rm Sub}^{p}_{G}$, then $T={\rm Id}$.
Proof.
Since $T$ acts trivially on $G/L$ and $L$, we have that $T$ acts distally on $G$. In fact, all the eigenvalues of ${\rm d\,}T$ are equal to 1 and $T$ is unipotent. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Let $U$ be a neighbourhood of the identity $e$ such that every element of $U$ is contained in a one-parameter subgroup. Since $G$ is connected, $U$ generates $G$ and it is enough to show that $T(x)=x$ for all $x\in U$.
Step 1: If possible, suppose $T(x)\neq x$ for some $x\in U$. There exists a one-parameter subgroup $\{x_{t}\}$ in $G$ such that $x_{1}=x$. Let $H=\overline{\{x_{t}\}L}$. Then $H$ is a closed connected solvable $T$-invariant subgroup of $G$ and $H/L$ is abelian. Let $\pi:G\to G/L$ be the natural projection. Then $\pi(H)=\overline{\pi(\{x_{t}\})}$. If $\pi(\{x_{t}\})$ is unbounded, it is closed and it is isomorphic to ${\mathbb{R}}$ and in this case $\{x_{t}\}$ is also unbounded and closed. Moreover $H=\{x_{t}\}L$.
Now suppose $\pi(\{x_{t}\})$ is relatively compact. As $H$ is solvable, there exists a maximal compact (connected, abelian) subgroup $K$ in $H$ such that $H=KL$. As $T(x)\neq x$ and $T|_{L}={\rm Id}$, we have that $T(h)\neq h$ for some $h\in K$, and hence $T(h)\neq h$ for some finite order element $h$ in $K\cap U$. There exists a closed (compact) one-parameter subgroup $\{h_{t}\}$ in $K$ such that $h_{1}=h$. Replacing $\{x_{t}\}$ by $\{h_{t}\}$, we have in this case too that $H$ is closed and $H=\{x_{t}\}L$.
Now we have that $H=\{x_{t}\}L$ is closed, where either $\pi(\{x_{t}\})$ (as well as $\{x_{t}\}$) is isomorphic to ${\mathbb{R}}$ or $\{x_{t}\}$ is isomorphic to ${\mathbb{S}}^{1}$ and $T(x_{1})\neq x_{1}$. As $T$ acts trivially on $G/L$, we have that $T(H)=H$ and $T(x_{t})=x_{t}y_{t}$, for some $y_{t}\in L$, $t\in{\mathbb{R}}$, where $y_{1}\neq e$.
Step 2: Let $A:=\{x_{t}\}$ and let $\{n_{k}\}\subset{\mathbb{N}}$ be an unbounded sequence such that $T^{n_{k}}(A)\to B$ for some $B\in{\rm Sub}^{p}_{H}$. By Lemma 2.2, $B$ is not discrete as $A$ is connected and non-trivial. Let $b\in B$ be such that $b\neq e$. There exist $t_{k}\in{\mathbb{R}}$, $k\in{\mathbb{N}}$, such that
$T^{n_{k}}(x_{t_{k}})=x_{t_{k}}y_{t_{k}}^{n_{k}}\to b.$
As $y_{t}\in L$, $t\in{\mathbb{R}}$, we get that $\pi(x_{t_{k}})\to\pi(b)$. If $H/L$ is isomorphic to ${\mathbb{R}}$, then we get that the sequence $\{t_{k}\}$ is bounded. In the second case where $\{x_{t}\}$ is compact, we can choose $\{t_{k}\}$ to be bounded. Passing to a subsequence if necessary, we may assume that $t_{k}\to t_{0}$. Therefore, $x_{t_{k}}\to x_{t_{0}}$ and $y_{t_{k}}\to y_{t_{0}}$ and $b=x_{t_{0}}y^{\prime}$, where $y_{t_{k}}^{n_{k}}\to y^{\prime}$ in $L$.
Step 3: We first assume that $L$ is simply connected. We show that $y_{t_{0}}=e$ in this case. Let ${\mathfrak{l}}$ be the Lie algebra of $L$. Since $L$ is simply connected and nilpotent, $\exp:{\mathfrak{l}}\to L$ is a homeomorphism with $\log$ as its inverse. Let $y_{t_{k}}=\exp Y_{k}$ for some $Y_{k}\in{\mathfrak{l}}$. Then $y_{t_{k}}^{n_{k}}=\exp(n_{k}Y_{k})$ and as $y_{t_{k}}^{n_{k}}\to y^{\prime}$, we have that $n_{k}Y_{k}\to\log y^{\prime}$ in ${\mathfrak{l}}$. As ${\mathfrak{l}}$ is isomorphic to ${\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$, it follows that $Y_{k}\to 0$ in ${\mathfrak{l}}$, and hence $y_{t_{k}}\to e$, i.e. $y_{t_{0}}=e$. Now
$T(b)=T(x_{t_{0}})y^{\prime}=x_{t_{0}}y_{t_{0}}y^{\prime}=x_{t_{0}}y^{\prime}=b.$
Thus $T(B)=B$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(A)=A$. Since $A$ is a one-parameter subgroup in $G$ and $T$ is unipotent, we get that $T(x_{t})=x_{t}$, $t\in{\mathbb{R}}$. This leads to a contradiction as $T(x_{1})\neq x_{1}$. Therefore, $T(x)=x$ for all $x\in U$, and hence, for all $x\in G$. Thus $T={\rm Id}$ if $L$ is simply connected.
Step 4: Suppose that $L$ is not simply connected. Let $M$ be the maximal compact (normal) subgroup of $L$. Then $M$ is central in $G$ and $L/M$ is simply connected. Let $\varrho:G\to G/M$ be the natural projection and let $\overline{T}$ be the automorphism on $G/M$ induced by $T$. Then $\overline{T}$ acts trivially on both $\varrho(L)$ and $\varrho(G)/\varrho(L)$. By Proposition 3.5, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$. From Step 3, we get that $\overline{T}$ acts trivially on $G/M$, i.e. $T(g)\in gM$, $g\in G$. In particular, for $\{x_{t}\}$ as above, $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in M$, for all $t\in{\mathbb{R}}$, and $y_{1}\neq e$. Since $M$ is central in $G$, $y_{t}x_{s}=x_{s}y_{t}$, $t,s\in{\mathbb{R}}$, and $\{y_{t}\}$ is a one-parameter subgroup of the compact central subgroup $M$. Note that $\{y_{t}\}$ is closed if $\{x_{t}\}$ is compact.
As ${\rm Sub}^{p}_{G}$ is compact, there exists an unbounded sequence $\{n_{k}\}$ in ${\mathbb{N}}$ such that $T^{n_{k}}(\{x_{t}\})\to C$ for some $C\in{\rm Sub}^{p}_{G}$. We show that $C=\{x_{t}\}\times\overline{\{y_{t}\}}$. As $\{x_{t}\}\subset\{x_{t}\}\times\overline{\{y_{t}\}}$, and the latter is $T$-invariant, we get that $C\subset\{x_{t}\}\times\overline{\{y_{t}\}}$. For a fixed $t\in{\mathbb{R}}$,
$T^{n_{k}}(x_{t/n_{k}})=x_{t/n_{k}}y_{t/n_{k}}^{n_{k}}=x_{t/n_{k}}y_{t}\to y_{t}$
as $n_{k}\to\infty$. Therefore, $y_{t}\in C$ for each $t$, and $\overline{\{y_{t}\}}\subset C$ as $C$ is closed. Also, $T^{n_{k}}(x_{t})=x_{t}y_{t}^{n_{k}}\to x_{t}y$ along a subsequence of $\{n_{k}\}$ for some $y\in\overline{\{y_{t}\}}\subset C$. Hence, $x_{t}\in C$, $t\in{\mathbb{R}}$. Thus $\{x_{t}\}\times\overline{\{y_{t}\}}\subset C$. Now $T(C)=C$, and we get that $T^{n_{k}}(T(\{x_{t}\}))\to C$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(\{x_{t}\})=\{x_{t}\}$, and since $T$ is unipotent, $T(x_{t})=x_{t}$ for each $t$. But $T(x_{1})=x_{1}y_{1}\neq x_{1}$. This leads to a contradiction. Therefore, $T(x)=x$ for each $x\in U$, and hence, for each $x\in G$. Thus $T={\rm Id}$. ∎
Proof of Theorem 1.3.
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. Suppose $T$ is unipotent and it acts distally on ${\rm Sub}^{p}_{G}$. We show that $T={\rm Id}$.
Step 1: Suppose $G$ is nilpotent. If possible, suppose $T\neq{\rm Id}$. Since $T$ is unipotent, by Proposition 3.10 of [42], there exists an increasing sequence of closed connected normal T-invariant subgroups $\{e\}=G_{0}\subset G_{1}\subset\cdots\subset G_{n}=G$, $n\geq 2$, such that $T$ acts trivially on each successive quotient group $G_{i}/G_{i-1}$, $1\leq i\leq n$, and $T$ does not act trivially on $G_{i+1}/G_{i-1}$, $11\leq i\leq n-$ 1. Then $G_{2}\neq G_{1}$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, it acts distally on ${\rm Sub}^{p}_{G_{2}}$. Now by Proposition 4.2, $T|_{G_{2}}={\rm Id}$, which leads to a contradiction. Therefore, $T={\rm Id}$ when $G$ is nilpotent.
Step 2: Suppose $G$ is not nilpotent. Let $N$ be the nilradical of $G$. As $T$ acts distally on ${\rm Sub}^{p}_{N}$, it follows from Step 1 that $T|_{N}={\rm Id}$. Suppose $G$ is solvable. As noted in Remark 3.3, we have that the automorphism induced by $T$ on $G/N$ is a finite order automorphism, and it is trivial since it is unipotent. Now by Proposition 4.2, $T={\rm Id}$.
Step 3: Suppose $G$ is not solvable. Let $R$ be the radical of $G$. Then as $T$ acts distally on ${\rm Sub}^{p}_{R}$, we have from Step 2 that $T|_{R}={\rm Id}$.
Now $G/N$ is reductive. Let $\pi:G\to G/N$ be the natural projection. Then $\pi(G)=SR^{\prime}$, an almost direct product of a semisimple Levi subgroup $S$ of $\pi(G)$ and the radical $R^{\prime}=\pi(R)$ of $\pi(G)$, where $R^{\prime}$ is central in $\pi(G)$. Let ${\overline{T}}$ be the automorphism of $\pi(G)$ induced by $T$. Then ${\overline{T}}$ acts trivially on $R^{\prime}$. Also ${\overline{T}}$ keeps $S$ invariant as it is the commutator subgroup of $\pi(G)$. Now if ${\overline{T}}$ acts trivially on $\pi(G)=G/N$, then by Proposition 4.2, $T={\rm Id}$. If possible suppose ${\overline{T}}$ does not act trivially on $\pi(G)$. As ${\overline{T}}$ is unipotent, there exists a nontrivial unipotent element $u$ in $S$ such that ${\overline{T}}={\rm inn}(u)$. Considering the Iwasawa decomposition $S=KAU$, where $u\in U$, we have that $u$ is not in the center of $AU$ (see Lemma 6.50 of [29]). Now $AUR^{\prime}$ is connected, solvable and ${\overline{T}}$ invariant. Let $B$ be the closure of $\pi^{-1}(AUR^{\prime})$. It is a closed connected solvable $T$-invariant subgroup of $G$. As $T$ acts distally on ${\rm Sub}^{p}_{B}$, by Step 2, $T|_{B}={\rm Id}$. This implies that ${\overline{T}}$ acts trivially on $AU$ and that $u$ centralises $AU$, a contradiction. Thus ${\overline{T}}$ acts trivially on $\pi(G)=G/N$, and as noted above, by Proposition 4.2, we get that $T={\rm Id}$.
The converse statement in the theorem is obvious. ∎
Proof of Theorem 1.1$(2)$ for non-nilpotent case.
Let $G$ be a connected Lie group in class ${\mathcal{C}^{\prime}}$. Then ${\rm Aut}(G)$ is almost algebraic as a subgroup of ${\rm GL}({\mathfrak{g}})$ (via the correspondence $T\leftrightarrow{\rm d\,}T$) (cf. [12]). Suppose $T\in{\rm Aut}(G)$ acts distally on ${\rm Sub}^{p}_{G}$. Suppose $G$ is not nilpotent. In particular, it is not a vector group. By Theorem 1.2, $T$ acts distally on $G$. Then for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected abelian almost algebraic subgroup of ${\rm Aut}(G)$ and it has the form ${\mathcal{K}}\times\,{\mathcal{U}}$, where ${\mathcal{K}},\,{\mathcal{U}}\subset{\rm Aut}(G)$, ${\mathcal{K}}$ is an abelian compact connected group and ${\mathcal{U}}$ is an abelian unipotent connected group; i.e. it consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now $T^{n}=S\circ\psi=\psi\circ S$, where $S\in{\mathcal{U}}$ is unipotent and $\psi\in{\mathcal{K}}$. Hence, $S$ acts distally on ${\rm Sub}^{p}_{G}$ (cf. [41], Lemma 2.2). By Theorem 1.3, $S={\rm Id}$. Thus $T^{n}=\psi\in{\mathcal{K}}$, and hence $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse is obvious (cf. [31], Lemma 4.2). ∎
Now we want to prove Theorem 1.4. We need an elementary result about certain automorphisms keeping a maximal compact connected abelian subgroup (maximal torus) invariant. In a connected Lie group $G$, all maximal tori are conjugate to each other. Any compact abelian subgroup of $G$ is contained in a maximal torus, as any maximal compact abelian subgroup of $G$ is connected. If $A$ is a maximal torus in $G$, then so is $T(A)$, for any $T\in{\rm Aut}(G)$, and $T(A)$ is conjugate to $A$. Any maximal compact subgroup of a connected solvable Lie group is abelian (cf. [25]).
In [15], any subgroup of ${\rm Aut}(G)$ which fixes elements of a maximal torus is shown to be almost algebraic. Given any automorphism $T$ of $G$, it need not keep any torus invariant. However, this is the case if $T$ is contained in a compact connected abelian subgroup of ${\rm Aut}(G)$, as shown in the following elementary lemma. The lemma also holds if we replace the radical $R$ by any closed connected characteristic subgroup of $G$.
Lemma 4.3.
Let $G$ be a connected Lie group. If ${\mathcal{K}}$ is a compact connected abelian subgroup of ${\rm Aut}(G)$, then there exists a maximal torus of $G$ (resp. of the radical of $G$) on which every automorphism in ${\mathcal{K}}$ acts trivially.
Proof.
Let $A$ be any maximal torus in a connected Lie group $G$. It is easy to see that for any closed connected normal subgroup $L$ of $G$, $A\cap L$ is a maximal torus of $L$.
Let ${\mathcal{K}}$ be a compact connected abelian subgroup of ${\rm Aut}(G)$ and let $H={\mathcal{K}}\ltimes G$; where the action of ${\mathcal{K}}$ on $G$ is by automorphisms. Then $H$ is a connected Lie group and $G$ is normal in $H$. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$ and hence normal in $H$. Let $A$ be a maximal torus in $H$ containing ${\mathcal{K}}$. Arguing as above, we get that $A\cap G$ (resp.  $A\cap R$) is a maximal torus in $G$ (resp. $R$). As ${\mathcal{K}}\subset A$ and $A$ is abelian, we get that every automorphism in ${\mathcal{K}}$ acts trivially on $A\cap G$ (resp. $A\cap R$). ∎
Proof of Theorem 1.4.
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. The statements $(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious. Now it is enough to show that $(1)\implies(5)$ when $G$ is not a vector group and $T$ acts distally on the largest compact connected central subgroup (say) $M$ of $G$. Suppose (1) holds, i.e. $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. If $M$ is trivial, then $G$ belongs to the class ${\mathcal{C}^{\prime}}$ and the assertion follows from Theorem 1.1 (2). Now suppose $M$ is non-trivial. By Theorem 1.2, $T$ acts distally on $G$. Then by Theorem 3.1 of [34], $\overline{T}$ acts distally on $G/M$, where $\overline{T}$ is the automorphism of $G/M$ induced by $T$.
Suppose $G=M$. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], $T^{n}$ is unipotent for some $n\in{\mathbb{N}}$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, so does $T^{n}$, and by Theorem 1.3, $T^{n}={\rm Id}$. Hence the assertion holds when $G$ is compact and abelian.
Now suppose $M$ is non-trivial and $G\neq M$. We show that there exists a maximal torus of $G$ which is $T^{k}$-invariant for some $k\in{\mathbb{N}}$. Let $\pi:G\to G/M$ be the natural projection. By Proposition 3.5, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$.
Suppose $G/M$ is a vector group. Then by Theorem 1.1 (1), $\overline{T}=\psi\circ\,\alpha\,{\rm Id}$, for some $\alpha\in{\mathbb{R}}\setminus\{0\}$ and $\psi$ generates a relatively compact subgroup in ${\rm Aut}(G/M)$. Since $\overline{T}$ acts distally on $G/M$, we get that $|\alpha|=1$, and hence $\overline{T}$ is contained in a compact subgroup of ${\rm Aut}(G/M)$. Now suppose $G/M$ is not a vector group. As $G/M$ belongs to the class ${\mathcal{C}^{\prime}}$, by Theorem 1.1 (2), $\overline{T}$ is contained in a compact subgroup of ${\rm Aut}(G/M)$. So in either case we have that $\overline{T}$ is contained in a compact subgroup of ${\rm Aut}(G/M)$.
Let ${\mathcal{K}}$ be the connected component of the identity in the closed (compact) subgroup generated by $\overline{T}$ in ${\rm Aut}(G/M)$. Then $(\overline{T})^{k}\in{\mathcal{K}}$ for some $k\in{\mathbb{N}}$. Now by Lemma 4.3, $(\overline{T})^{k}$ keeps a maximal torus (say) $A$ of $G/M$ invariant and it acts trivially on $A$. Let $K=\pi^{-1}(A)$. Then $K$ is a maximal torus of $G$ and $T^{k}(K)=K$.
Note that $K$ is compact, connected and abelian. Moreover, as $T$ acts distally on $G$, so does $T^{k}$ and the restriction of $T^{k}$ to $K$ is also distal. Arguing as above for $K$ instead of $M$, we get that for some $m\in{\mathbb{N}}$, $T^{m}|_{K}$ is unipotent and, as it acts distally on ${\rm Sub}^{p}_{K}$, it acts trivially on $K$ (by Theorem 1.3). Without loss of any generality, we may replace $T$ by $T^{m}$ and assume that $T$ fixes every element of $K$, i.e. $T\in F_{K}(G):=\{\varrho\in{\rm Aut}(G)\mid\varrho(x)=x\mbox{ for all }x\in K\}$. We know that $F_{K}(G)$ is almost algebraic (cf. [15]). As $T$ acts distally on $G$ and $T\in F_{K}(G)$, for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathcal{B}}\times{\mathcal{U}}$ of $F_{K}(G)$, where ${\mathcal{B}}$ is a compact group and ${\mathcal{U}}$ consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now we have that $T=\phi\circ S=S\circ\phi$ such that $\phi\in{\mathcal{B}}$ and $S\in{\mathcal{U}}$. As ${\mathcal{B}}$ is compact, we have that $S$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.3, $S={\rm Id}$. Then $T=\phi\in{\mathcal{B}}$ and (5) holds. ∎
Using Theorems 1.1 and 1.4 and some known results, we prove Theorem 1.5.
Proof of Theorem 1.5.
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Suppose ${\mathcal{H}}$ acts distally on the largest compact connected central subgroup of $G$. Then so does $\overline{{\mathcal{H}}}$. Note that ${\mathcal{H}}$ acts distally on a compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$ (both the above statements follow from Theorem 1 in [16]). Also (7) holds for ${\mathcal{H}}$ if and only if it holds when ${\mathcal{H}}$ is replaced by $\overline{{\mathcal{H}}}$, as ${\mathcal{K}}{\mathcal{D}}$ is closed in ${\rm GL}(d,{\mathbb{R}})$ for ${\mathcal{K}},{\mathcal{D}}$ as in (7). Therefore, to prove the assertions, we may assume that ${\mathcal{H}}$ is closed.
Note that every compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [41], Lemma 2.2). Now $(6)\implies(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious.
First suppose that $G$ is not a vector group. We show that $(1)\implies(6)$. Suppose (1) holds, i.e. every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.4, every $T\in{\mathcal{H}}$ is contained in a compact subgroup of ${\rm Aut}(G)$. Recall that ${\rm Aut}(G)$ is identified with a closed subgroup of ${\rm GL}({\mathfrak{g}})$ via the map $T\mapsto{\rm d\,}T$, where ${\mathfrak{g}}$ is the Lie algebra of $G$ and it is a $d$-dimensional vector space for some $d\in{\mathbb{N}}$ (cf. [9], see also [22]). Therefore, ${\mathcal{H}}$ is also a closed subgroup of ${\rm GL}({\mathfrak{g}})$ and the latter is isomorphic to ${\rm GL}(d,{\mathbb{R}})$, and we get by Theorem 1.1 of [17] (see also Proposition 2 of [39]) that ${\mathcal{H}}$ is compact. Thus (6) holds and $(1-6)$ are equivalent. Thus, if $G$ is not a vector group, the assertions in the theorem hold for this case.
Now suppose $G={\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$. Suppose (3) holds, i.e. ${\mathcal{H}}$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$. Then $T\in\rm{(NC)}$ for every $T\in{\mathcal{H}}$. Since $G$ belongs to the class ${\mathcal{C}^{\prime}}$, by Theorem 4.1 of [42], $T$ is contained in a compact subgroup of ${\rm Aut}(G)$ for every $T\in{\mathcal{H}}$. Since ${\mathcal{H}}$ is closed, arguing as above using Theorem 1.1 of [17] or Proposition 2 of [39], we get that ${\mathcal{H}}$ is compact. Hence (6) holds. Thus $(3-6)$ are equivalent when $G$ is a vector group. Now we show that (1), (2) and (7) are equivalent. Suppose (7) holds. Note that ${\mathcal{D}}=\{r\,{\rm Id}\in{\rm GL}(d,{\mathbb{R}})\mid r\in{\mathbb{R}}\setminus\{0\}\}$. Therefore, ${\mathcal{D}}$ acts trivially on ${\rm Sub}^{p}_{G}$ and the action of ${\mathcal{K}}{\mathcal{D}}$ on ${\rm Sub}^{p}_{G}$ is same as that of the compact group ${\mathcal{K}}$ on ${\rm Sub}^{p}_{G}$. Therefore, the action of ${\mathcal{K}}{\mathcal{D}}$, and hence, the action of ${\mathcal{H}}$ on ${\rm Sub}^{p}_{G}$ is distal. Thus $(7)\implies(2)$. We have already noted that $(2)\implies(1)$.
Now we show that $(1)\implies(7)$. Suppose every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. Let $T\in{\mathcal{H}}$ be fixed. By Theorem 1.1 (1), $T\in{\mathcal{B}}{\mathcal{D}}$, for some compact group ${\mathcal{B}}\subset{\rm GL}(d,{\mathbb{R}})$. Let ${\mathcal{H}}_{1}:=\overline{[{\mathcal{H}},{\mathcal{H}}]}$. Then ${\mathcal{H}}_{1}\subset{\mathcal{H}}$ as the latter is closed. Now let $T\in{\mathcal{H}}_{1}$. Then $\det T=1$. From above we have that $T=\phi\,\circ\,r\,{\rm Id}$ for some $\phi\in{\mathcal{B}}$ and some $r\in{\mathbb{R}}\setminus\{0\}$. Since ${\mathcal{B}}$ is a compact group, $\det\phi=\pm 1$, and hence $r=\pm 1$. Therefore, $T$ generates a relatively compact group, for every $T\in{\mathcal{H}}_{1}$. As ${\mathcal{H}}_{1}$ is a closed subgroup of ${\rm GL}(d,{\mathbb{R}})$ by Theorem 1.1 of [17] or Proposition 2 of [39], ${\mathcal{H}}_{1}$ is compact. Let $N({\mathcal{H}}_{1})$ be the normaliser of ${\mathcal{H}}_{1}$ in ${\rm GL}(d,{\mathbb{R}})$. Then $N({\mathcal{H}}_{1})$ is a closed algebraic subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{H}},{\mathcal{D}}\subset N({\mathcal{H}}_{1})$. Moreover, $N({\mathcal{H}}_{1})\subset N({\mathcal{H}}_{1}{\mathcal{D}})$. Being algebraic, $N({\mathcal{H}}_{1})$ has finitely many connected components, and hence so does $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$. In particular any closed abelian subgroup of $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ is compactly generated (cf. [24]). Let $\rho:N({\mathcal{H}}_{1})\to N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ be the natural projection. Then $\rho({\mathcal{H}})$ is abelian, and hence so is $\overline{\rho({\mathcal{H}})}$. Therefore, $\overline{\rho({\mathcal{H}})}$ is compactly generated, and it has a unique maximal compact subgroup, say $\mathcal{L}$.
[…]
Now we state two corollaries which are easy to deduce using results proven above and a result in [1].
Corollary 4.4.
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Then ${\mathcal{H}}$ acts distally on both $G$ and ${\rm Sub}^{p}_{G}$ if and only if $\overline{{\mathcal{H}}}$ is compact.
Corollary 4.4 is a direct consequence of Theorem 1.5 in case $G$ is not a vector group. If $G$ is a vector group, then as ${\mathcal{H}}$ acts distally on $G$, the eigenvalues of every element in ${\mathcal{H}}$ have absolute value 1 (cf. [1]), and this, together with the statement (7) in Theorem 1.5, is equivalent to the statement that $\overline{{\mathcal{H}}}$ is compact. We omit the details.
The following corollary is an extension of Corollary 4.6 of [42] for a particular class of Lie groups. It can be proven easily using Theorems 1.1 and 1.5 and Proposition 3.5. We omit the proof.
Corollary 4.5.
Let $G$ be a connected Lie group such that $G/M$ is not a vector group, where $M$ is the maximal compact connected central subgroup of $G$. Let $T\in{\rm Aut}(G)$ (resp. ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$). If $T$ (resp. ${\mathcal{H}}$) acts distally on ${\rm Sub}^{p}_{G}$, then $T$ (resp. ${\mathcal{H}}$) has bounded orbits; i.e. $\{T^{n}(x)\mid n\in{\mathbb{Z}}\}$ (resp. ${\mathcal{H}}x$) is relatively compact in $G$, for every $x\in G$.
Now we prove Corollary 1.6. The statement in the corollary which requires a proof is $(1)\implies(5)$ and can be proven along the lines of proof of $(1)\implies(5)$ of Corollary 4.5 in [42]; instead of Corollary 3.9 and Theorem 4.3 there, one has to use Theorems 1.2 and 1.3. We will give a proof here for the sake of completeness.
Proof of Corollary 1.6.
Let $G$ be a connected Lie group. The statement in (5) implies that ${\rm Inn}(G)$ is compact, and hence it follows that $(5)\implies(4)$. Also $(4)\implies(3)\implies(1)$ and $(4)\implies(2)\implies(1)$ are obvious. It is enough to show that $(1)\implies(5)$.
Suppose (1) holds; i.e. every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$. If $G$ is either compact or a vector group, then (5) holds trivially. Now suppose $G$ neither compact nor a vector group. Since every inner automorphism acts trivially on the center of $G$, we have by Theorem 1.2 that it acts distally on $G$. Therefore, $G$ is distal (cf. [36]), i.e. the action of ${\rm Inn}(G)$ on $G$ is distal. It follows from Theorem 9 of [36] and Corollary 2.1 (ii) of [27] that $G/R$ is compact, where $R$ is the radical of $G$. Let $N$ be the nilradical of $G$ and let $x\in N$. Then ${\rm inn}(x)$ is unipotent, and by Theorem 1.3, ${\rm inn}(x)={\rm Id}$. This implies that $x$ belongs to the center of $G$ for every $x\in N$. Thus $N$ is abelian and central in $G$. Now for the radical $R$, we know that $[R,R]\subset N$ is central in $G$, and hence $R$ is nilpotent and it is same as $N$ and it is abelian and central in $G$. Now $G/N$ is compact, where $N$ is central in $G$, i.e. $N={\mathbb{R}}^{n}\times M$, for some $n\in{\mathbb{N}}$, where $M$ is compact. Here, $V:={\mathbb{R}}^{n}$ is a vector subgroup which is central. By Lemma 3.7 of [25], $G={\mathbb{R}}^{n}\times K$, where $K$ is the maximal compact subgroup of $G$. Thus (5) holds. ∎
The distal actions of automorphisms $T$ of a connected Lie group $G$ on ${\rm Sub}^{a}_{G}$ were studied earlier if $G$ is in class ${\mathcal{C}^{\prime}}$ or $T$ is unipotent. Here we have studied a more general case when $T$ acts distally on the maximal central torus of $G$ and the action is considered on ${\rm Sub}^{p}_{G}$, a smaller subspace of ${\rm Sub}^{a}_{G}$. More generally, we have also studied the actions of subgroups ${\mathcal{H}}$ of ${\rm Aut}(G)$ on ${\rm Sub}^{p}_{G}$ under a condition that the ${\mathcal{H}}$-action on the maximal central torus of $G$ is distal. This latter condition is satisfied if ${\mathcal{H}}$ is contained in ${\rm Aut}(G)^{0}$ or more generally, if it is contained in an almost algebraic subgroup of ${\rm Aut}(G)$.
The following example of a connected solvable Lie group $G$, known as the Walnut group, is not covered by Theorem 1.1 or any earlier theorems for the distal action of a general automorphism on ${\rm Sub}^{a}_{G}$, as it has nontrivial central torus and ${\rm Aut}(G)$ is not almost algebraic (see [15]). But it is covered by theorems proven here.
Example 4.6.
Let $\mathbb{H}$ be the 3-dimensional Heisenberg group consisting of  $\,3\times 3$ real strictly upper triangular matrices. There is a canonical action of $SL(2,{\mathbb{R}})$ on $\mathbb{H}$ which fixes elements of its center, which is isomorphic to ${\mathbb{R}}$. Let $N=\mathbb{H}/D$, where $D$ is a discrete central subgroup isomorphic to ${\mathbb{Z}}$. Then $N$ is a connected 2-step nilpotent group and the ${\rm SL}(2,{\mathbb{R}})$-action on $\mathbb{H}$ carries over to the action on $N$. Let $G=SO(2,{\mathbb{R}})\ltimes N$, the Walnut group. Since the center of $G$ is isomorphic to ${\mathbb{S}}^{1}$, ${\rm Aut}(G)$ acts distally on it. Thus Theorems 1.2 and 1.4 hold for the action of any $T\in{\rm Aut}(G)$ and Theorem 1.5 holds for any subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$.
One can also take a non-solvable group $G=SL(2,{\mathbb{R}})\ltimes N$, for $N$ mentioned as above, whose automorphism group is almost algebraic (see [15]). Theorems mentioned in the above example hold also for this group. This group $G$ is mentioned in Remark 4.8 of [42], although it has a maximal torus of dimension 2, ${\rm Aut}(G)$ is almost algebraic (see [15]), and the conclusion in Theorem 4.1 of [42] mentioned in the remark holds for this $G$.
It would be interesting to characterise general automorphisms of any connected Lie group $G$ which act distally on ${\rm Sub}_{G}$ or its closed invariant subspace without any condition involving the action on the central torus.
Acknowledgement. The authors would like to acknowledge the support in part by the International Centre for Theoretical Sciences (ICTS) during a visit for participating in the program - ICTS Ergodic Theory and Dynamical Systems (code: ICTS/etds-2022/12).